# Probing Cosmic Curvature with Fast Radio Bursts and DESI DR2

### v1

@@ Abstract
The spatial curvature of the Universe remains a central question in modern cosmology. In this work, we explore the potential of localized Fast Radio Bursts (FRBs) as a novel tool to constrain the cosmic curvature parameter $\Omega_{k}$ in a cosmological model-independent way. Using a sample of 80 FRBs with known redshifts and dispersion measures, we reconstruct the Hubble parameter $H(z)$ via artificial neural networks, which is then used to obtain angular diameter distances $D_{A}(z)$ through two complementary approaches. First, we derive the comoving distance $D_{C}(z)$ and $D_{A}(z)$ from FRBs without assuming a fiducial cosmological model. Then, the $H(z)$ reconstruction is combined with Baryon Acoustic Oscillation (BAO) measurements to infer $D_{A}(z)$. By comparing the FRB-derived and BAO+FRB-derived $D_{A}(z)$ values, we extract constraints on $\Omega_{k}$ under the Friedmann–Lemaître–Robertson–Walker metric. Our results show consistency with a spatially flat Universe within $1\sigma$ uncertainties, although a mild preference for negative curvature is observed across both covariance-based and Gaussian analyses. This study highlights the growing relevance of FRBs in precision cosmology and demonstrates their synergy with BAO data as a powerful probe of the large-scale geometry of the Universe.
@@ I Introduction
The current constraints on $\Omega_{k}$ are largely consistent with flatness, although uncertainties remain. For instance, combining Planck Cosmic Microwave Background (CMB) lensing data with low-redshift Baryon Acoustic Oscillations (BAO) measurements yields $\Omega_{k}=0.0007\pm 0.0019$ Aghanim et al. (2020). However, other analyses challenge this result. Notably, Di Valentino et al. (2020) reported a preference for a mildly closed Universe, with $\Omega_{k}=-0.044^{+0.018}_{-0.015}$, based solely on CMB temperature and polarization data from Planck satellite. Another study found a preference for a closed Universe at the $2\sigma$ confidence level, measuring $\Omega_{k}=-0.089^{+0.049}_{-0.046}$ when applying Effective Field Theories of Large Scale Structure (EFTofLSS) to BAO data Glanville et al. (2022). On the other hand, Wu and Zhang (2024) reported $\Omega_{k}=0.108\pm 0.056$, which is in $2.6\sigma$ tension with the CMB lensing data result. To obtain this estimate, they used a combination of datasets, including BAO, Cosmic Chronometers (CC), and Strong Gravitational Lensing (SGL).
However, it is important to recognize that these constraints are inherently cosmological model-dependent, relying on assumptions within the non-flat $\Lambda$CDM framework, which suffers from the well-known intrinsic degeneracy between cosmic curvature and the Hubble constant. Moreover, as with the Hubble tension, the apparent disagreement in curvature estimates raises the question of whether systematic biases or new physics are at play. In this context, exploring methods that provide a complementary and cosmological model-independent perspective is particularly valuable. Developing curvature measurements that are independent of the cosmological model, especially those that do not rely solely on CMB-based assumptions, offers a promising pathway to gaining a better understanding of these issues.
@@
Beyond their intriguing astrophysical nature, FRBs have proven to be valuable probes for cosmology. In particular, localized FRBs, those with measured redshifts and host galaxy associated, provide unique observational access to the otherwise elusive ionized content of the intergalactic medium (IGM). Over the past few years, multiple studies have employed these events to estimate key cosmological quantities, such as the Hubble constant $H_{0}$ Macquart et al. (2020); Wu et al. (2022); Zhao et al. (2022); Kalita et al. (2025); Hagstotz et al. (2022); Fortunato et al. (2025), cosmographic parameters Fortunato et al. (2023), the cosmic proper distance Yu and Wang (2017), the baryon fraction in the IGM Li et al. (2020); Yang et al. (2022); Lemos et al. (2022), constrain dark energy equation of state Zhou et al. (2014); Gao et al. (2014), trace the magnetic fields in the IGM Akahori et al. (2016), probe the equivalence principle Wei et al. (2015); Tingay and Kaplan (2016); Nusser (2016), the rest mass of photons Wu et al. (2016); Shao and Zhang (2017); Lin et al. (2023), to cite some.
@@ II Basic properties of Fast Radio Bursts
[…]
where $c$ is the speed of light, $G$ the gravitational constant, and $m_{p}$ the proton mass. $\Omega_{b}$ is the cosmic baryon density, and $E(z)=H(z)/H_{0}$ the normalized Hubble parameter. The term $f_{\mathrm{IGM}}(z)$ denotes the baryon fraction in the intergalactic medium, which we take as $f_{\mathrm{IGM}}=0.82\pm 0.04$ Shull et al. (2012); Zhou et al. (2014). The factor $f_{e}(z)$ describes the baryons ionization state and is defined as:
@@ III Distances in cosmology
A related quantity is the horizon distance $D_{H}$, which represents the maximum scale over which information could have been exchanged since the Big Bang. It is defined as:
$D_{H}=\frac{c}{H_{0}}.$ (8)
This distance sets a fundamental scale in cosmology and is often used to normalize other cosmological distances.
@@
$D_{M}=\begin{cases}\frac{D_{H}}{\sqrt{\Omega_{k}}}\sinh\left(\sqrt{\Omega_{k}}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}>0,\\ D_{C},&\text{if }\Omega_{k}=0,\\ \frac{D_{H}}{\sqrt{|\Omega_{k}|}}\sin\left(\sqrt{|\Omega_{k}|}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}<0,\end{cases}$ (9)
@@ IV.1 Extracting $D_{C}$ and $D_{A}$ from $\langle\mathrm{DM_{IGM}}\rangle$
[…]
@@ IV.2 Estimation of the cosmic curvature
  • Covariance matrix approach: In our first approach, we account for the statistical correlation among the distance data points, primarily because both the angular diameter distance $D_{A}(z)$ and the comoving distance $D_{C}(z)$ depend on the same reconstructed Hubble parameter $H(z)$. To properly account for these correlations in the likelihood analysis, we construct a covariance matrix that incorporates both statistical uncertainties and correlated terms, following the strategy proposed in Liu et al. (2025). The likelihood function is given by
@@
    The statistical part captures the propagated uncertainties from the reconstruction of $H(z)$ and the BAO-based measurements. The correlated term reflects the intrinsic coupling between $D_{A}$ and $D_{C}$, since both quantities are derived from the same FRB-based estimates of $H(z)$. This term is computed using the standard definition of covariance: $\mathrm{Cov}(X_{i},Y_{j})=\mathbb{E}[(X_{i}-\mu_{i})(Y_{j}-\mu_{j})]$.
    Although the DESI collaboration provides a fiducial covariance matrix, it is derived under specific model assumptions and tailored to their reconstruction method. In contrast, our analysis relies on a model-independent reconstruction of $H(z)$ and therefore requires a covariance matrix that accurately reflects our methodology. By constructing the covariance matrix directly from our reconstructed distances, we ensure internal consistency and avoid potential mismatches. Moreover, the authors in Liu et al. (2025) have shown that this method yields results consistent with the DESI covariance estimates, further supporting the reliability of this approach.
@@
    where $\sigma_{i}$ represents the combined error from the reconstruction of $D_{A}$ and $D_{C}$. This method does not account for correlations among the data points and may underestimate the uncertainty, but it serves as a useful baseline.
@@ V Data
To implement our analysis, we use a subset of 80 from the 92 localized FRBs compiled by Wang et al. (2025). In addition to requiring known redshifts and dispersion measures, we impose a selection criterion also introduced in Wang et al. (2025) to ensure the physical consistency of the intergalactic medium contribution to the dispersion measure, $\mathrm{DM}_{\mathrm{IGM}}$. Specifically, we require
@@
Data source $z$ $D_{M}/r_{d}$ $D_{H}/r_{d}$ Ref.
BOSS galaxy–galaxy 0.38 $10.23\pm 0.17$ $25.00\pm 0.76$ Alam et al. (2017)
eBOSS galaxy–galaxy 0.51 $13.36\pm 0.21$ $22.33\pm 0.58$ Alam et al. (2017)
eBOSS galaxy–galaxy 0.70 $17.86\pm 0.33$ $19.33\pm 0.62$ Bautista et al. (2021); Gil-Marín et al. (2020)
eBOSS galaxy–galaxy 0.85 $19.60\pm 0.41$ $19.60\pm 0.72$ Tamone et al. (2020); De Mattia et al. (2021)
eBOSS galaxy–galaxy 1.48 $30.21\pm 0.79$ $13.23\pm 0.47$ Neveux et al. (2020); Hou et al. (2021)
DESI DR2 LRG1 0.510 $13.588\pm 0.167$ $21.863\pm 0.425$ Karim et al. (2025)
DESI DR2 LRG2 0.706 $17.351\pm 0.177$ $19.455\pm 0.330$ Karim et al. (2025)
DESI DR2 LRG3+ELG1 0.934 $21.576\pm 0.152$ $17.641\pm 0.193$ Karim et al. (2025)
DESI DR2 ELG2 1.321 $27.601\pm 0.318$ $14.176\pm 0.221$ Karim et al. (2025)
DESI DR2 QSO 1.484 $30.512\pm 0.760$ $12.817\pm 0.516$ Karim et al. (2025)
Table 1: Measurements of $D_{M}/r_{d}$ and $D_{H}/r_{d}$ from various surveys.
In addition, seven events were excluded from our analysis for specific reasons. First, FRB200110E is omitted because it is located at an extremely close distance of only $3.6~{}\mathrm{Mpc}$, within a globular cluster in the M81 galaxy Kirsten et al. (2022); Bhardwaj et al. (2021). Its proximity implies a negligible contribution from the intergalactic medium, making it unsuitable for cosmological inference. Second, FRB210117 and FRB190520 display dispersion measures significantly exceeding theoretical expectations based on their redshift. These anomalies classify them as statistical outliers, probably influenced by a dense or highly magnetized local environment in their host galaxy Bhandari et al. (2023); Gordon et al. (2023); Niu et al. (2022). Moreover, we did not reconstruct the functions at higher redshift values due to the lack of FRB data points in this regime. Although there are four points around $z\sim 1$ — FRB220610, FRB221029, FRB230521, FRB240123 — we excluded them because they could bias the reconstruction, given the gap between them and the last point used in this study at $z=0.66$. The remaining 80 FRBs form a clean and physically consistent dataset for the curvature analysis presented in the following sections.

### v2

@@ Abstract
The spatial curvature of the Universe remains a central question in modern cosmology. In this work, we explore the potential of localized Fast Radio Bursts (FRBs) as a novel tool to constrain the cosmic curvature parameter $\Omega_{k}$ in a cosmological model–independent way. Using a sample of 120 FRBs with known redshifts and dispersion measures, we reconstruct the Hubble parameter $H(z)$ via artificial neural networks, and use it to obtain angular–diameter distances $D_{A}(z)$ through two complementary approaches. First, we derive the comoving distance $D_{C}(z)$ and $D_{A}(z)$ directly from FRBs without assuming a fiducial cosmology. Then, we combine the FRB-based $H(z)$ with Baryon Acoustic Oscillation (BAO) DESI DR2 measurements to infer $D_{A}(z)$. By comparing the FRB-derived and BAO+FRB-derived $D_{A}(z)$, we constrain spatial curvature. Our covariance-based likelihood (accounting for correlated uncertainties) yields $\Omega_{k}=-0.31\pm 0.57$, while a diagonal (Gaussian) treatment gives $\Omega_{k}=-0.13\pm 0.46$. Both estimations are consistent with spatial flatness at the $1\sigma$ level, albeit with a mild preference for negative curvature. Explicitly accounting for the full covariance broadens the intervals and avoids underestimation of uncertainties. These results highlight the growing relevance of FRBs in precision cosmology and their synergy with BAO as a powerful, cosmological model–independent probe of the large–scale geometry of the Universe.
@@ I Introduction
The current constraints on $\Omega_{k}$ are largely consistent with flatness, although uncertainties remain. For instance, combining Planck Cosmic Microwave Background (CMB) lensing data with low-redshift Baryon Acoustic Oscillations (BAO) measurements yields $\Omega_{k}=0.0007\pm 0.0019$ Aghanim et al. (2020). However, other analyses challenge this result. Notably, Di Valentino et al. (2020) reported a preference for a mildly closed Universe, with $\Omega_{k}=-0.044^{+0.018}_{-0.015}$, based solely on CMB temperature and polarization data from Planck satellite. Another study found a preference for a closed Universe at the $2\sigma$ confidence level, measuring $\Omega_{k}=-0.089^{+0.049}_{-0.046}$ when applying Effective Field Theories of Large Scale Structure (EFTofLSS) to BAO data Glanville et al. (2022). On the other hand, Wu and Zhang (2024) reported $\Omega_{k}=0.108\pm 0.056$, which is in $2.6\sigma$ tension with the CMB lensing data result. To obtain this estimate, they used a combination of datasets, including BAO, Cosmic Chronometers (CC), and Strong Gravitational Lensing (SGL). Other recent studies reached conclusions consistent with spatial flatness by combining alternative datasets, reporting mean values of $\Omega_{k}$ that deviate slightly from zero but remain within $1\sigma$ of flatness. For instance, in Vagnozzi et al. (2021a) used full-shape galaxy power spectra from BOSS DR12 CMASS and Planck data, finding $\Omega_{k}=0.0023\pm 0.0028$, while Vagnozzi et al. (2021b) combined Planck with CC data to obtain $\Omega_{k}=-0.0054\pm 0.0055$.
[…]
@@
Beyond their intriguing astrophysical nature, FRBs have proven to be valuable probes for cosmology. In particular, localized FRBs, those with measured redshifts and host galaxy associated, provide unique observational access to the otherwise elusive ionized content of the intergalactic medium (IGM). Over the past few years, multiple studies have employed these events to estimate key cosmological quantities, such as the Hubble constant $H_{0}$ Macquart et al. (2020); Wu et al. (2022); Zhao et al. (2022); Kalita et al. (2025); Hagstotz et al. (2022); Fortunato et al. (2025), cosmographic parameters Fortunato et al. (2023), the cosmic proper distance Yu and Wang (2017), the baryon fraction in the IGM Li et al. (2020); Yang et al. (2022); Lemos et al. (2022), constrain dark energy equation of state Zhou et al. (2014a); Gao et al. (2014), trace the magnetic fields in the IGM Akahori et al. (2016), probe the equivalence principle Wei et al. (2015); Tingay and Kaplan (2016); Nusser (2016), the rest mass of photons Wu et al. (2016); Shao and Zhang (2017); Lin et al. (2023), to cite some.
@@
Throughout this work, the expression model-independent denotes that no early-universe calibration or specific cosmological model is assumed: BAO enter only through dimensionless ratios whose combination cancels $r_{d}$, and $H(z)$ is reconstructed directly from FRB data via the Macquart relation. This approach relies on Eqs. (4) and (5) (defined in Sec. II), which connect the FRB dispersion measures to $H(z)$ via minimal physical assumptions (FLRW geometry and standard ionization fractions) without assuming a particular cosmological model. Curvature effects enter explicitly only in the FRB-only through the $S_{k}$ function mapping from $D_{C}$ to $D_{M}$.
@@ II Basic properties of Fast Radio Bursts
$\langle\mathrm{DM}_{\mathrm{IGM}}\rangle=\left(\frac{3c}{8\pi Gm_{p}}\right)\Omega_{b}H_{0}\int_{0}^{z}\frac{(1+z^{\prime})f_{\mathrm{IGM}}(z^{\prime})f_{e}(z^{\prime})}{E(z^{\prime})}dz^{\prime},$ (4)
where $c$ is the speed of light, $G$ the gravitational constant, and $m_{p}$ the proton mass. $\Omega_{b}$ is the cosmic baryon density, and $E(z)=H(z)/H_{0}$ the normalized Hubble parameter. The term $f_{\mathrm{IGM}}(z)$ denotes the baryon fraction in the intergalactic medium, which we take as $f_{\mathrm{IGM}}=0.82\pm 0.04$ Shull et al. (2012); Zhou et al. (2014b). The factor $f_{e}(z)$ describes the baryons ionization state and is defined as:
@@ III Distances in cosmology
A related scale is the Hubble distance (or Hubble radius) at the present epoch,
$D_{H_{0}}=\frac{c}{H_{0}},$ (8)
with the general redshift-dependent definition
$D_{H}(z)=\frac{c}{H(z)}.$ (9)
This scale is often used to normalize cosmological distances and should not be confused with causal horizons (particle or event horizons), which involve time integrals of $H(z)$.
@@
$D_{M}=\begin{cases}\frac{D_{H_{0}}}{\sqrt{\Omega_{k}}}\sinh\left(\sqrt{\Omega_{k}}\frac{D_{C}}{D_{H_{0}}}\right),&\text{if }\Omega_{k}>0,\\ D_{C},&\text{if }\Omega_{k}=0,\\ \frac{D_{H_{0}}}{\sqrt{|\Omega_{k}|}}\sin\left(\sqrt{|\Omega_{k}|}\frac{D_{C}}{D_{H_{0}}}\right),&\text{if }\Omega_{k}<0,\end{cases}$ (10)
@@
The equation above provides an angular diameter distance estimate that combines BAO ratios with an FRB-based reconstruction of the Hubble parameter. In this construction, the curvature parameter $\Omega_{k}$ does not enter explicitly, since the BAO measurements appear only through the ratio of their reported quantities and the sound horizon cancels. By contrast, the FRB-only approach reconstructs the comoving distance from the dispersion measure and then maps it into a transverse comoving distance using Eq. (9), which explicitly involves $\Omega_{k}$. Thus, curvature dependence arises only in the FRB-only estimator, while the BAO+FRB estimator remains algebraically independent of $\Omega_{k}$.
@@ IV.1 Extracting $D_{C}$ and $D_{A}$ from $\langle\mathrm{DM_{IGM}}\rangle$
$\langle\mathrm{DM_{IGM}}(z)\rangle=A\int_{0}^{z}F(z^{\prime})\,\dot{D}_{C}(z^{\prime})dz^{\prime},$ (15)
@@ IV.2 Estimation of the cosmic curvature
Covariance matrix approach: In our first approach, we account for the statistical correlation among the distance data points, primarily because both the angular diameter distance $D_{A}(z)$ and the comoving distance $D_{C}(z)$ depend on the same reconstructed Hubble parameter $H(z)$. To properly account for these correlations in the likelihood analysis, we construct a covariance matrix that incorporates both statistical uncertainties and correlated terms, following the strategy found in Liu et al. (2025), based on the error propagation theory Owen (2000). The likelihood function is given by
@@
The DESI DR2 survey Karim et al. (2025) provides the covariance of the primary BAO observables $\mathbf{x}\equiv\big(D_{M}/r_{d},\;D_{H}/r_{d}\big)$ at each redshift, with a full (off–diagonal) covariance matrix $\mathbf{C}_{x}$. Since our analysis uses Eq. (14), we propagate BAO errors to $D_{A}$ via the Jacobian Owen (2000) $\mathbf{J}_{i}=\big(\partial f/\partial(D_{M}/r_{d}),\;\partial f/\partial(D_{H}/r_{d})\big)\big|_{z_{i}}$:
$\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}=\mathbf{J}\,\mathbf{C}_{x}\,\mathbf{J}^{\!\top}.$ (21)
Following Eq. (20), only its diagonal enters $\mathbf{Cov}^{\mathrm{stat}}$, and we add in quadrature the FRB-inferred $H(z)$ contribution on the data side:
$\big(\mathbf{Cov}^{\mathrm{stat}}\big)_{ii}=\big(\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}\big)_{ii}+\left(\frac{\partial D_{A}}{\partial H}\sigma_{H}\right)^{2}_{z_{i}}.$ (22)
The correlated term captures the intrinsic coupling between $D_{A}^{\mathrm{FRB}}$ and $D_{C}^{\mathrm{FRB}}$ induced by the same FRB–reconstructed $H(z)$. We estimate it empirically from the FRB ensembles (bootstrap realizations) as a cross–covariance:
$\mathbf{Cov}^{\mathrm{corr}}_{ij}$ $=$ $\mathrm{Cov}\!\left[D_{A}^{\mathrm{FRB}}(z_{i}),\,D_{C}^{\mathrm{FRB}}(z_{j})\right]$ (23)
$=$ $\mathbb{E}\!\left[\big(D_{A}^{\mathrm{FRB}}-\mu_{A}\big)_{i}\,\big(D_{C}^{\mathrm{FRB}}-\mu_{C}\big)_{j}\right].$ (24)
Although the DESI collaboration provides a fiducial covariance matrix, it is derived under specific model assumptions and tailored to their reconstruction method. In contrast, our analysis relies on a model-independent reconstruction of $H(z)$ and therefore requires a covariance matrix that accurately reflects our methodology. By constructing the covariance matrix directly from our reconstructed distances, we ensure internal consistency and avoid potential mismatches. As a cross-check, we repeated the inference using the DESI-provided covariance matrix (projected in the same way) and obtained posteriors for $\Omega_{k}$ that were statistically indistinguishable from our baseline – central values and uncertainties changed negligibly.
@@
where $\sigma_{i}$ represents the combined error from the reconstruction of $D_{A}$ and $D_{C}$. This method does not account for correlations among the data points and may underestimate the uncertainty, but it serves as a useful baseline. The likelihood therefore compares a curvature–independent quantity, defined in Eq. (13), with a curvature–dependent one obtained via Eq. (9), and the sensitivity to $\Omega_{k}$ emerges solely from this comparison.
In summary, the COV analysis uses the full matrix $\mathbf{Cov}=\mathrm{diag}\!\big(\mathbf{Cov}^{\mathrm{stat}}\big)+\mathbf{Cov}^{\mathrm{corr}}$, where the BAO contribution to $D_{A}$ is rigorously propagated from DESI’s (off–diagonal) covariance via the Jacobian, while the correlated FRB term is measured from the FRB ensemble. For comparison, the Gaussian case adopts only the diagonal variance $\sigma_{i}^{2}=\big(\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}\big)_{ii}+\left(\partial D_{A}/\partial H\,\sigma_{H}\right)^{2}_{z_{i}}$, i.e. it neglects $\mathbf{Cov}^{\mathrm{corr}}$.
@@ V Data
To implement our analysis, we use a subset of 120 from the 131 localized FRBs presented in table LABEL:tab:frb_sample. In addition to requiring known redshifts and dispersion measures, we impose a selection criterion also introduced in Wang et al. (2025) to ensure the physical consistency of the intergalactic medium contribution to the dispersion measure, $\mathrm{DM}_{\mathrm{IGM}}$. Specifically, we require
@@
Tracer $z$ $D_{M}/r_{d}$ $D_{H}/r_{d}$ Ref.
LRG1 0.510 $13.588\pm 0.167$ $21.863\pm 0.425$ Karim et al. (2025)
LRG2 0.706 $17.351\pm 0.177$ $19.455\pm 0.330$ Karim et al. (2025)
LRG3+ELG1 0.934 $21.576\pm 0.152$ $17.641\pm 0.193$ Karim et al. (2025)
ELG2 1.321 $27.601\pm 0.318$ $14.176\pm 0.221$ Karim et al. (2025)
QSO 1.484 $30.512\pm 0.760$ $12.817\pm 0.516$ Karim et al. (2025)
Lya 2.330 $338.988\pm 0.531$ $8.632\pm 0.101$ Karim et al. (2025)
Table 1: Measurements of $D_{M}/r_{d}$ and $D_{H}/r_{d}$ from DESI DR2 survey.
In addition, seven events were excluded from our analysis for specific reasons. First, FRB200110E is omitted because it is located at an extremely close distance of only $3.6~\mathrm{Mpc}$, within a globular cluster in the M81 galaxy Kirsten et al. (2022); Bhardwaj et al. (2021). Its proximity implies a negligible contribution from the intergalactic medium, making it unsuitable for cosmological inference. Second, FRB210117 and FRB190520 display dispersion measures significantly exceeding theoretical expectations based on their redshift. These anomalies classify them as statistical outliers, probably influenced by a dense or highly magnetized local environment in their host galaxy Bhandari et al. (2023); Gordon et al. (2023); Niu et al. (2022). A practical limitation of our reconstruction is the sparse sampling at high redshift ($z\gtrsim 1$). Our baseline analysis includes all available localized FRBs after cleaning the dataset (120 events in total), among which only a handful lie near or above $z\sim 1$ (notably FRB220610, FRB221029, FRB230521, FRB240123, and the most distant FRB240304B at $z=2.148$). In this regime, the ANN is weakly constrained by the data, which naturally broadens the predictive variance and, consequently, inflates the propagated uncertainties in $H(z)$, $D_{C}(z)$, and $D_{A}^{\mathrm{FRB}}(z)$. This effect is expected and reflects limited leverage rather than a modeling bias. Since our curvature inference is most strongly informed by the redshift range where BAO measurements exist ($z\lesssim 2.5$), the net impact is a conservative widening of error bands at high $z$ without qualitatively altering the central trends used to constrain $\Omega_{k}$.